OpenStudy (anonymous):

Professor Mean takes a 4-sided die to the grocery store, starts at one end of the chips aisle, and travels to the other end. At each different kind of chips, Dr. Mean rolls the die. If it comes up a 4, he purchases the chips for the party. There are 40 different kinds of chips in the aisle. Define random variable x = number of types of chips purchased out of 40. What is the mean of the random variable x? What is the standard deviation of the random variable x?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

What have you got so far? Or were you just saying hi to the stats crowd? :D

OpenStudy (anonymous):

i think i know what to do

OpenStudy (anonymous):

the "there are 40 different kins of chips" thing is what makes me think im doing it wrong

OpenStudy (anonymous):

i was reading it what did it say?

OpenStudy (anonymous):

It was just the formula for expected value, but I deleted it because I misstyped it. I will be back in a few minutes, sorrry.

OpenStudy (misssunshinexxoxo):

Professor Mean? woah

OpenStudy (misssunshinexxoxo):

Devation is 6 because there are only 6 different numerical otcomes

OpenStudy (misssunshinexxoxo):

@xartaan I got this

OpenStudy (anonymous):

sry i was afk xD

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did i lose connection?

OpenStudy (ybarrap):

It's 4-sided. Is each side equally likely to occur?

OpenStudy (anonymous):

yea

OpenStudy (ybarrap):

So, what is the chance that a 4 will show if you tossed the die once, if each side is equally likely?

OpenStudy (anonymous):

1/4

OpenStudy (ybarrap):

So, now 1 out of every 4 tosses will be (on average) a 4. You agree?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

Ok, if I tossed the die 4 times how many times would 4 show up (on average)?

OpenStudy (anonymous):

once

OpenStudy (ybarrap):

Correct! If I tossed the die 8 times, how many times would 4 show up (on average)?

OpenStudy (anonymous):

twice

OpenStudy (ybarrap):

That's right! So for every 4 tosses, I get 1 4 (on average). So if I toss it 40 times, how many times does 4 show up (on average)?

OpenStudy (anonymous):

6 times

OpenStudy (ybarrap):

When we tossed it 4 times We get a 4 (on average) \(4 \times\cfrac{1}{4}=1\) fours _______________________________________________________ When we tossed it 8 times We get a 4 (on average) \(8 \times\cfrac{1}{4}=2\) fours _______________________________________________________ Do you see the pattern?

OpenStudy (ybarrap):

If we tossed it 40 times, We get a 4 (on average) \(40\times\cfrac{1}{4}=?\) fours

OpenStudy (anonymous):

hmm i see its 8

OpenStudy (anonymous):

so hen what is the mean of variable x then is it the same thing

OpenStudy (anonymous):

40 x 1/2?

OpenStudy (ybarrap):

(making sure you are checking) The mean number value of x is the average number of times that 4 shows up when you through the die 40 times. $$ 40\times\cfrac{1}{4} $$

OpenStudy (ybarrap):

*throw

OpenStudy (anonymous):

okay i got hat i was just confused with the 1/2

OpenStudy (anonymous):

so is the standard dev is 6 like that other person said?

OpenStudy (anonymous):

Not quite,, recall the definition of the Var is E[X^2] - [E(X)^2] .. You know the mean is 10 so that's the easy half... Sorry trying to do this from my phone, hope that's clear this far though

OpenStudy (ybarrap):

I assume that you understand now how to get the average N * p, where p = 1/4. Next, the standard deviation is computed similarly $$ \large{ \sigma^2=Npq } $$ Where q=1-p So the standard deviation is $$ \sigma=\sqrt{40\times\cfrac{1}{4}\times\cfrac{3}{4}}\approx2.74 $$ These formulas assume that this whole process can be modeled by the binomial distribution - http://en.wikipedia.org/wiki/Binomial_distribution. It is binomial because the die either IS or IS NOT a 4 - and so you can think of this as either being a 1 or a 0 -- binary and hence the binomial distribution. I hope this made sense.